Circular No. 8666

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

(617) PATROCLUS F. Marchis, J. Berthier, D. Hestroffer, P. Descamps, and the Keck

Science team report on a campaign of observations of the (617) Patroclus system performed using the Keck 10-m-telescope Laser Guide Star Adaptive Optics system in 2004–2005. The observations indicate that the two components, separated by 680 ± 20 km, revolve around their center of mass in 4.283 ± 0.004 days in a roughly circular orbit with a pole at ecliptic coordinates λ = 236^{o} ± 5^{o}, β = −61^{o} ± 1^{o} (equinox J2000.0). The solution, obtained over four years of observations, including also Gemini Science Archive data, is purely Keplerian. Considering thermal measurements taken in Nov. 2000 by Fernandez et al. (2003, A.J. 126, 1563), and assuming the same composition for the spherical components (R_{1} = 61 km and R_{2} = 56 km), the bulk density derived is quite low (0.8 ± 0.2 g/cm^{3}; Marchis et al. 2006, Nature 439, 565-567). The IAU Committee on Small-Body Nomenclature has approved the permanent designation and name (617) Patroclus I = Menoetius (father of Patroclus) for the system’s smaller component, which was designated S/2001 (617) 1 when discovered in Sept. 2001 by W. J. Merline et al. (IAUC 7741).

SUPERNOVA 2006T IN NGC 3054 L. A. G. Monard, Pretoria, South Africa, reports his discovery of an

apparent supernova (mag ∼ 17.2 ± 0.4) on unfiltered CCD images taken with a 0.30-m reflector on Jan. 30.988 UT, with confirming images on Jan. 31.847 showing the new object at mag ∼ 17.4 ± 0.2. SN 2006T is located at α = 9^{h}54^{m}30^{s}.21, δ = −25^{o}42′29′′.3 (equinox 2000.0; USNO-A2.0 reference stars), which is 22′′ east and 21′′ south of the nucleus of NGC 3054. Nothing is visible at this location on a Digitized Sky Survey image (limiting red mag 20.5) or on an image taken on 2006 Jan. 16.955 by Monard (limiting red mag 18.0). Note that SN 2004de also occurred in this galaxy.

SUPERNOVA 2006E IN NGC 5338 H. Yamaoka, Kyushu University, writes that K. Itagaki (Yamagata,

Japan, 0.30-m reflector) found a prediscovery image of SN 2006E (cf. IAUC 8658, 8660) at red mag ∼ 13.6 on an unfiltered CCD exposure of NGC 5338 taken on Jan. 2.835 UT.

2006 February 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green